An orifice meter has an orifice of 10 cm diameter and a coefficient of discharge of 0.65. If it fixed in a pipe of 25 cm diameter with flowing oil of 0.8 specific gravity. The pressure difference between the pipe and the orifice is measured by a mercury manometer that gives a reading of

80 cm. determine the flow rate.


1
Expert's answer
2021-07-19T16:03:38-0400

Gives

"d_0=10cm"

Area "a_0=\frac{\pi}{4}\times10^2=78.5cm^2"

Daimeter of pipe "d_1=25cm"

Area

"a_1=\frac{\pi}{4}\times25^2=490.625cm^2"

"s_0=0.8"

Reading of different manometer x=80cm of Mercury


"h=x [\frac{S_g}{S_0}-1]=50\times [\frac{13.6}{0.8}-1]=800cm"

h=800cm Of oil

Rate of flow

"Q=c_d \frac{a_0a_1\times \sqrt{2gh}}{\sqrt {a_1^2-a_0^2}}"


"Q=0.65\times\frac{78.5\times490.625\times \sqrt{2\times981\times490.625}}{\sqrt {490.625^2-78.5^2}}"

"Q=\frac{0.65\times78.5\times490.625\times981.124}{484.30}=50715.66cm^3/sec"